OpenStudy (anonymous):

A car traveling at 10.0 m/s accelerates at the rate of 0.8 m/s2 for an interval of 2 s. Find the final velocity.

OpenStudy (anonymous):

v=u+at: v=10+0.8*2; =11.6 m/s

OpenStudy (anonymous):

Thanks so much!! I wasn't taught this equation so this makes a huge difference. Although, if you don't mind me asking, what is u? Is it just the unknown variable?

OpenStudy (anonymous):

u=inital velocity a=acceleration t=time period

OpenStudy (anonymous):

v=final velocity

OpenStudy (anonymous):

Okie dokies! Thanks so very much!

OpenStudy (anonymous):

your welcome...